write an equation for the following conic. if it is a parabola, it has a vertex at the origin, and if it is an ellipse or a hyperbola, it is centered at the origin. focus at (4,0) and e= 1/2
Answer by lwsshak3(11628)

write an equation for the following conic. if it is a parabola, it has a vertex at the origin, and if it is an ellipse or a hyperbola, it is centered at the origin. focus at (4,0) and e= 1/2
**
if 0 (x-h)^2/a^2+(y-k)^2/b^2=1, a>b, (h,k)=(x,y) coordinates of center
center: (0,0) (given)
c=4
c^2=16
e=1/2=c/a
a=2c=8
a^2=64
c^2=a^2-b^2
b^2=a^2-c^2=64-16=48
Equation of ellipse:
x^2/64+y^2/48=1